18 July 2017
Tuesday 18th of July 2017

Wataru and Jan in the undisturbed grand white landscape 20 km SE of EGRIP camp.

Drilling again

In the afternoon the bore hole was finally determined to be free of slush after retrieving more than 200 kg during the past days filtering operation. Hence drilling commenced again with two runs of good quality cores before dinner. Logging the cores without introducing breaks is, however, becoming increasingly difficult.

During the afternoon Wataru and Jan made radar measurements 20 km SE of camp, crossing the shear margin on skidoo without problems. They also had a bit of time to enjoy the peaceful grand white landscape with camp completely out of sight.

Back in camp, the AWI drill system was put on an air force pallet and the packing and weighing of ice core boxes as well as other cargo continued. Scientific programs on the surface and in the science trench continued as well, albeit with Ina as the only scientist in the physical properties laboratory today.

Chris took care to groom the skiway after yesterday’s snowfall, and in anticipation of the upcoming flight period.

What we did today:

  1. Finished filtering of bore hole, more than 200 kg of slush retrieved in total.
  2. Deep drilling core length at 19.00: 2.99 m.
  3. Logging depth at 19.00: 672.31 m.
  4. Radar measurements 20km SE from camp.
  5. AWI drill system palletized.
  6. Documentation and strapping of ice core boxes.
  7. Water vapour, methane, aerosols, snow and boundary layer measurements
  8. Physical properties measurements.
  9. Water isotope sampling for sublimation studies.
  10. Finished mounting end stops on core troughs.
  11. Grooming skiway.

Weather: Mostly sunny. Wind jumped from NE to SSW during the morning, up to 5 m/s. Temperatures dropping from -11°C to -17°C during the day.

FL, Bo Vinther

← Previous entry   Next entry →